OpenStudy (anonymous):

How do you graph f(x) = x-3

OpenStudy (anonymous):

will you be using a graphing calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok. so the equation f(x)=x-3 also means y=x-3. now we need to make a table of x and y. we'll choose integers for x and plug them in to obtain a value for y.

OpenStudy (anonymous):

so, lets start with 0. When x = 0. we get y = (0) - 3, which gives us y = -3

OpenStudy (anonymous):

so on the graph, at x = 0, y = -3. so put a point on coordinate (0,-3). and you continue to x = 1, x = 2, x =3 for as many points as you want. also do this for negative x values. x = -1, x = -2. you'll soon realize that this is a linear function (straight line). so therefore you don't need to continue forever you just extrapolate the line to fit the graph.

OpenStudy (anonymous):

wow this is godsend thank you

OpenStudy (anonymous):

Draw a coordinate plane. Draw the line y = x i.e. a line with slope 1 through the origin (0,0) Move the line from the point (0,0) where the line y = x intersects down to point (0,-3) (translation of the graph)

OpenStudy (anonymous):

thank you for your time I can't believe I didn't get this the first time the examples in my math book are not nearly as understandable as you guys